Opened 10 years ago

Closed 9 years ago

Last modified 2 years ago

#234 closed flight processing (fixed)

GB08/15, flight day 078b/2009, Eaves Wood

Reported by: jatt Owned by: knpa
Priority: alpha 4 low Milestone: 2009 Data processing completion
Component: Archiving Keywords:
Cc: Other processors: mark1

Description (last modified by knpa)

Data location: ~arsf/arsf_data/in_progress/2009/flight_data/uk/GB08_15-2009_078b_Eaves_Wood

Data arrived from ARSF via SATA disk N on 30th March 2009.

Scientific details: /users/rsg/arsf/arsf_data/2008/ARSF_Applications/GB_and_NL/GB08-15 Science Case.pdf

Priority: a4L

PI: Dr Alan Blackburn

Sensors:

  • Eagle (18/12/2009)
  • Hawk (18/12/2009)
  • ALS-II Lidar (??/11/2009)
  • Photography (06/01/2010)

Quicklook

TBD.

Attachments (1)

GB08_15rollVsTime.png (38.0 KB) - added by knpa 10 years ago.

Download all attachments as: .zip

Change History (37)

comment:1 Changed 10 years ago by jatt

  • Owner mggr deleted

comment:2 Changed 10 years ago by mggr

  • Type changed from task to flight processing

comment:3 Changed 10 years ago by knpa

  • Priority changed from alpha 4 high to alpha 4 low

comment:4 Changed 10 years ago by knpa

  • Owner set to knpa
  • Status changed from new to assigned

comment:5 Changed 10 years ago by knpa

  • Owner knpa deleted
  • Status changed from assigned to new

Advised that the boresight for this project is variable as it is pre day-130.

comment:6 Changed 10 years ago by mggr

  • Component changed from ARSF to Processing: general

comment:7 Changed 10 years ago by knpa

  • Owner set to knpa

comment:8 Changed 10 years ago by knpa

started processing of the lidar (with variable boresight)

Changed 10 years ago by knpa

comment:9 Changed 10 years ago by knpa

Lidar processing completed. Delivery procedure underway.

comment:10 Changed 10 years ago by knpa

Delivery procedure completed (still needs to be verified by someone).
Waiting for vectors, have e-mailed PI asking if they would like a pre-vector delivery.

comment:11 Changed 10 years ago by knpa

  • Owner changed from knpa to mark1

PI has confirmed they would like pre-vector delivery on Lidar.
Await delivery check and then will completed delivery.

comment:12 Changed 10 years ago by knpa

Logsheet gave an incorrect project code (GB08_02), this has been edited in the copy sent in the delivery.

comment:13 Changed 10 years ago by knpa

DELIVERED

comment:14 Changed 10 years ago by mark1

  • Owner mark1 deleted

comment:15 Changed 10 years ago by knpa

delivery comment above refering to LIDAR, eagle and hawk still to process

comment:16 Changed 10 years ago by mark1

  • Owner set to mark1
  • Status changed from new to assigned

comment:17 Changed 10 years ago by knpa

  • Owner changed from mark1 to knpa

comment:18 Changed 10 years ago by mark1

  • Other processors set to mark1

comment:19 Changed 10 years ago by mark1

Have corrected project code on original logsheet also.

comment:20 Changed 10 years ago by mark1

Also renamed flight line numbers on logsheet from CROSS and CROSS to CROSS1 and CROSS2

comment:21 Changed 10 years ago by emca

Delivery check for Eagle & Hawk complete, all data present and correct. Ready to deliver.
(delivery directory located at ~airborne/workspace/GB08_15-2009_078b_Eaves_Wood/delivery/20091211)

comment:22 Changed 10 years ago by mark1

  • Description modified (diff)

Eagle and Hawk data sent on 18th December 2009.

comment:23 Changed 9 years ago by mark1

Digital Photography data dispatched to PI on 6th January 2010

comment:24 Changed 9 years ago by mark1

  • Description modified (diff)

comment:25 Changed 9 years ago by knpa

  • Component changed from Processing: general to Archiving

Archived by DTG (12/01/10).
Preparing for NEODC.

comment:26 Changed 9 years ago by knpa

  • Description modified (diff)

comment:27 Changed 9 years ago by knpa

Tarball created and NEODC informed ready for transfer.

comment:28 Changed 9 years ago by knpa

The tarball was ready to go to NEODC but it turns out photography was not rsynced before the archive process started. Mark1 has now rsynced this and it is ready to be submitted for archiving again.

comment:29 Changed 9 years ago by knpa

No rcd/logs or eventfile for this project.

No webcam for this project.

comment:30 Changed 9 years ago by mark1

Eagle Boresight GPT SCT
1 -0.04 -0.74 -0.25 0 0
2 -0.40 -0.58 0 0 0
3 -0.01 -0.43 0.39 0 0
4 -0.39 -0.72 -0.13 0 0
5 -0.01 -0.43 0.39 0 0
6 -0.32 -0.62 -0.13 0 0
7 -0.17 -0.48 0.40 0 0
8 -0.58 -0.73 0 0 0
9 0.17 -0.51 -0.61 0 0
10 -0.31 -0.69 -0.13 0 0

comment:31 Changed 9 years ago by mark1

Hawk Boresight GPT SCT
1 0.74 0.23 0.13 -0.12 0
2 -0.38 0.29 0.13 0 1.10
3 -0.76 0.18 0.12 0 0
4 -0.36 0.24 0.15 0 0
5 -0.59 0.29 0.13 0.08 0
6 -0.47 0.24 -0.10 0 0
7 -0.33 0.16 -0.14 0 0
8 -0.21 0.22 -0.19 0 0
9 -0.63 0.14 0.22 0.08 0

comment:32 Changed 9 years ago by mark1

No dark frames for hawk line 8 - used the ones from line 7

comment:33 Changed 9 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Archived and backed up by NEODC.
Closing ticket.

comment:34 Changed 6 years ago by tipo

Reprocessed nav and hyperspectral.

comment:35 Changed 3 years ago by lah

Archiving

Tidied and ready for upload to NEODC.

comment:36 Changed 2 years ago by asm

Archiving

Reprocessed hyperspectral has been archived and it is available from CEDA at ​http://browse.ceda.ac.uk/browse/neodc/arsf/2009/GB08_15/GB08_15-2009_078b_Eaves_Wood/hyperspectral-reprocessed
Plymouth 11th May 2017

Last edited 2 years ago by asm (previous) (diff)
Note: See TracTickets for help on using tickets.